# ISAC channel simulator GBSM

### v4

@@ Abstract
Integrated sensing and communication (ISAC) has been recognized as the key technology in the vision of the sixth generation (6G) era. With the emergence of new concepts in mobile communications, the channel model is the prerequisite for system design and performance evaluation. Currently, 3GPP Release 19 is advancing the standardization of ISAC channel models. Nevertheless, a unified modeling framework has yet to be established. This paper provides a simulation diagram of ISAC channel modeling extended based on the Geometry-Based Stochastic Model (GBSM), compatible with existing 5G channel models and the latest progress in the 3rd Generation Partnership Project (3GPP) standardization. We first introduce the progress of the ISAC channel model standardization in general. Then, a concatenated channel modeling approach is presented considering the team’s standardization proposals, which is implemented on the BUPTCMCC-6G-CMG+ channel model generator. We validated the model in cumulative probability density function (CDF) in statistical extension of angle and delay, and radar cross section (RCS). Simulation results show that the proposed model can realistically characterize the feature of channel concatenation and RCS within the ISAC channel.
@@
For ISAC channel simulation, a channel model that separately generates communication and sensing clusters based on the 3GPP model and utilizes sensing echoes for sensing has been proposed [19]. Some studies introduce the concept of shared clusters to explain the phenomenon of scattering body reuse between the communication and sensing channels [5, 20, 21]. Some studies focus on deterministic ISAC channel modeling [22, 23]. However, these studies do not align with the target and background channel modeling framework or the concatenated channel method in 3GPP Rel-19. Most of these approaches use Mono-static echoes for sensing and do not support Bi-static sensing modes.
@@
Objects creating hazards on road/railways | Highway, Urban grid, Urban grid HST |
@@
Different from existing ISAC channel simulation models [19, 20, 21, 22, 23], we propose a unified ISAC channel simulation framework based on the 3GPP Rel-19 ISAC channel fundamental framework, where the ISAC channel is the combination of the target and background channels. This framework maximizes the use and compatibility with existing 5G standardized channel models. This work integrates our team’s contributions to the standardization process, presenting a solution that aligns with the latest standardized progress and consensus in ISAC channel modeling [17].
@@ 2 Simulation Architecture
Figure 2 depicts the proposed simulation process extended by the 3GPP channel model[13, 14, 15]. In Step 1, users can perform general configurations, including antenna configuration, scenario settings, target types, and BS, UT, and target positions. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients. Steps 4, 9, and 12 correspond to large-scale parameter concatenation, RCS generation, and small-scale parameter concatenation. Step 15 refers to the combination of target and background channels.
@@ 2.1.1 Large scale parameters concatenation
where and represent the distances of the Tx-target and target-Rx, respectively. denotes the center frequency of the simulation. is the speed of light with m/s. is the mean RCS.
@@ 2.1.2 Small scale parameters concatenation
In this section, we derive and present a unified function for the ISAC target channel. This function is compatible with the concatenated channel methodology outlined in the 3GPP RAN1 ISAC channel modeling issue, incorporating key parameters such as RCS small-scale variations, RCS polarization, and Doppler. The unified function for each receiver and transmitter antenna element pair u and s in LOS+LOS condition is given in (2.1.2), while the equations for the remaining LOS+NLOS, NLOS+LOS, and NLOS+NLOS conditions can be obtained by simplifying 2.1.2.
@@
Here, and are the number of clusters in the Tx-target and target-Rx links, respectively. is the wavelength corresponding to the center carrier frequency. and are the number of rays in each cluster in the Tx-target and target-Rx links, respectively. and represent the field components of the receiving antenna in the zenith and azimuth directions, respectively. and represent the zenith and the azimuth angle, respectively. and represent the field components of the transmitting antenna in the zenith and azimuth directions, respectively. represents the RCS value related to angular terms, the detailed definition of which can be found in Section 2.2.2. and are the power of the -th cluster in the Tx-target link and the power of the -th cluster in the target-Rx link, respectively. and represent the position vectors of the receiving and transmitting antenna elements, respectively. and are the unit vectors in spherical coordinate, with the calculation expressions as follows:
@@
Where, and are the cross-polarization power ratio (XPR) in the Tx-target and target-Rx link, respectively. and represent the phase of the LOS ray of Tx-target and target-Rx links, respectively. Here, and are the three dimension (3D) distance of Tx-target and target-Rx links, respectively. S is the polarization matrix of the target RCS introduced in section 2.2.1.
@@ 2.2.1 RCS polarization
The scattering matrix S represents the coefficients related to the target’s scattering characteristics and polarization as follows:
@@ 2.2.2 RCS modeling based Measurement
where component is included in the large scale and is a deterministic value. The component is included in the small scale and is a deterministic angle-dependent value or scattering pattern based on measurements. The component is a normal distribution with a mean and variance. is used as in (1), while and are used as in (2.1.2), (2.1.2), (2.1.2), (6). Modeling the RCS of a UAV based on this framework, since its RCS is angle-dependent, the component is set to the mean or peak value, or simply set to 1, etc., the component is a function related to the angle, and is a random value independent of the angle, representing the fluctuation of RCS. This model framework can effectively model the RCS of sensing targets with different scattering patterns. In section 4, we present the measured results of the human target RCS and the simulation fitting results based on the proposed framework, which are used to comparatively validate the feasibility of the model.
@@ 3.1 Accuracy metric
Based on equation (22), we can drive . The angle resolution is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array . The basic principle of velocity measurement is based on the Doppler frequency. From Table 2, we obtain . Analogous to range resolution and angular resolution, Doppler resolution can be expressed as , in which is the number of transmitted pulses, and is the pulse repetition interval.
@@ 4.1 Concatenation methods
Metrics defined for the target channel include power, delay spread (DS), and angle spread (ASA, ASD, ZSA, ZSD). Based on the concatenation methods at the cluster level, ray level, random coupling, and order coupling, the 8 options above can be further subdivided into 10 cases:
@@ 5 Conclusion
[…]

### v5

# RESEARCH PAPER

@@ Abstract
Integrated sensing and communication (ISAC) has been recognized as the key technology in the vision of the sixth generation (6G) era. With the emergence of new concepts in mobile communications, the channel model is the prerequisite for system design and performance evaluation. Currently, the 3rd Generation Partnership Project (3GPP) Release 19 is advancing the standardization of ISAC channel models. Nevertheless, a unified modeling framework has yet to be established. This paper provides a simulation diagram of ISAC channel modeling extended based on the Geometry-Based Stochastic Model (GBSM), compatible with existing 5G channel models and the latest progress in the 3GPP Rel-19 standardization. We first introduce the progress of the ISAC channel model standardization in general. Then, a concatenated channel modeling method is presented considering the team’s standardization proposals, which is implemented on the BUPTCMCC-6G-CMG+ channel model generator. We validated the model in cumulative probability density function in statistical extension of angle and delay, and radar cross section (RCS). Simulation results show that the proposed model can realistically characterize the feature of channel concatenation and RCS within the ISAC channel.
@@
For ISAC channel simulation, a channel model that separately generates communication and sensing clusters based on the 3GPP model and utilizes sensing echoes for sensing has been proposed [19]. Some studies introduce the concept of shared clusters to explain the phenomenon of scattering body reuse between the communication and sensing channels [5, 20, 21, 22]. Some studies focus on deterministic ISAC channel modeling [23, 24]. However, these studies do not align with the target and background channel modeling framework or the concatenated channel method in 3GPP Rel-19. Most of these approaches use Mono-static echoes for sensing and do not support Bi-static sensing modes.
@@
| Objects creating hazards on road/railways | Highway, Urban grid, HST |
@@
Different from existing ISAC channel simulation models [19, 20, 21, 22, 23, 24], we propose a unified ISAC channel simulation framework based on the 3GPP Rel-19 ISAC channel fundamental framework, where the ISAC channel is the combination of the target and background channels. This framework maximizes the use and compatibility with existing 5G standardized channel models. This work integrates our team’s contributions to the standardization process, presenting a solution that aligns with the latest standardized progress and consensus in ISAC channel modeling [17].
@@ 2 Simulation Architecture
Figure 2 depicts the proposed simulation process extended by the 3GPP channel model[13, 14, 15]. In Step 1, users can perform general configurations, including antenna configuration, scenario settings, target types, and BS, UT, and target positions. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients. Steps 4, 9, and 12 correspond to large-scale parameter concatenation, RCS generation, and small-scale parameter concatenation, respectively. Step 15 refers to the combination of target and background channels.
@@ 2.1.1 Large scale parameters concatenation
where and represent the distances of the Tx-target and target-Rx, respectively. and represent the path loss of the Tx-target and target-Rx links, respectively. denotes the center frequency of the simulation. is the speed of light with m/s. is the mean RCS.
@@ 2.1.2 Small scale parameters concatenation
In this section, we derive and present a unified function for the ISAC target channel. This function is compatible with the concatenated channel methodology outlined in the 3GPP RAN1 ISAC channel modeling issue, incorporating key parameters such as RCS small-scale variations, target polarization, and Doppler. The unified function for each receiver and transmitter antenna element pair u and s in LOS+LOS condition is given in (2.1.2), while the equations for the remaining LOS+NLOS, NLOS+LOS, and NLOS+NLOS conditions can be obtained by simplifying 2.1.2.
@@
Here, and are the number of clusters in the Tx-target and target-Rx links, respectively. and represent the LOS ray’s delay of Tx-target and target-Rx links, respectively. and represent the delay of the -th ray within the -th cluster in the Tx-target link and the delay of the -th ray within the -th cluster in target-Rx link, respectively. Note that the delays referenced in this paper align with the relative delays specified in 3GPP TR 38.901 Section 7.5. For practical ISAC channel simulations, absolute delays need to be computed to evaluate sensing performance. For LOS condition, absolute delays can be calculated geometrically by , where is the 3D distance. For NLOS condition, Release 19 discussions are ongoing; meanwhile, Section 7.6.9 of TR 38.901 may serve as a provisional baseline. The actual propagation delay for ISAC channel can then be obtained by superimposing the absolute delay onto the relative delays generated in this work. is the wavelength corresponding to the center carrier frequency. and are the number of rays in each cluster in the Tx-target and target-Rx links, respectively. and represent the field components of the receiving antenna in the zenith and azimuth directions, respectively. and represent the zenith and the azimuth angle, respectively. and represent the field components of the transmitting antenna in the zenith and azimuth directions, respectively. represents the RCS value related to angular terms, the detailed definition of which can be found in Section 2.2.2. and are the power of the -th cluster in the Tx-target link and the power of the -th cluster in the target-Rx link, respectively. and represent the position vectors of the receiving and transmitting antenna elements, respectively. and are the unit vectors in spherical coordinate, with the calculation expressions as follows:
@@
Where, and are the cross-polarization power ratio (XPR) in the Tx-target and target-Rx link, respectively. and represent the phase of the LOS ray of Tx-target and target-Rx links, respectively. Here, and are the three dimension (3D) distance of Tx-target and target-Rx links, respectively. , , , and represent the polarization effects induced by the target for the LOS+LOS, LOS+NLOS, NLOS+LOS, and NLOS+NLOS link conditions, all of which can be calculated following (16).
@@ 2.2.1 RCS polarization
Here, and represent the vertical and horizontal polarization electric field component at the transmitter, respectively. and represent the vertical and horizontal polarization electric field component at the receiver, respectively. It should be noted that the electric field in (15) only considers the effect of polarization, and the modeling of RCS in this paper is independent of polarization. The target polarization matrix S is defined as follows:
@@ 2.2.2 RCS modeling based Measurement
where component is included in the large scale and is a deterministic value. The component is included in the small scale and is a deterministic angle-dependent value or scattering pattern based on measurements. The component is a log-normal distribution with a mean and variance. is used as in (1), while and are used as in (2.1.2), (2.1.2), (2.1.2), (6). Modeling the RCS of a UAV based on this framework, since its RCS is angle-dependent, the component is set to the mean or peak value, or simply set to 1, etc., the component is a function related to the angle, and is a random value independent of the angle, representing the fluctuation of RCS. This model framework can effectively model the RCS of sensing targets with different scattering patterns. In section 4, we present the measured results of the human target RCS and the simulation fitting results based on the proposed framework, which are used to comparatively validate the feasibility of the model.
@@ 3.1 Accuracy metric
Based on equation (22), we can drive . The angle resolution is defined as the minimum distinguishable angle, which is related to the main lobe width of the antenna array . The basic principle of velocity measurement is based on the Doppler frequency. From Table 2, we obtain . Analogous to range resolution , angular resolution and Doppler resolution can be expressed as , in which is the number of transmitted pulses, and is the pulse repetition interval.
@@ 4.1 Concatenation methods
Here, Case 1, 2, 3, and 4 represent link conditions such as LOS+LOS, LOS+NLOS, NLOS+LOS, and NLOS+NLOS respectively. Metrics defined for the target channel include power, delay spread (DS), and angle spread (ASA, ASD, ZSA, ZSD). Based on the concatenation methods at the cluster level, ray level, random coupling, and order coupling, the 8 options above can be further subdivided into 10 cases:
@@ 5 Conclusion
It is critical to emphasize that the Mono-static background channel cannot be directly modeled by repurposing existing channel modeling methodologies and parameters outlined in current 3GPP technical reports. RAN1 is actively progressing discussions on the development of Mono-static background channel modeling. Future research is required to address the modeling of the Mono-static background channel, as well as the spatial consistency modeling in the ISAC channel. Nevertheless, the proposed method for combining the target channel and background channel in the paper is universally applicable across various scenarios.
This work was supported in part by the National KeyR&D Program of China (Grant No. 2023YFB2904805), in part by the National Natural Science Foundation of China (Grant Nos. 62341128 and 62201087), in part by the Beijing Natural Science Foundation - Xiaomi Innovation Joint Fund (Grant No. L243002), in part by Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2023B0303000001) and in part by the Beijing University of Posts and Telecommunications-China Mobile Research Institute Joint Innovation Center.